### v2

# MERA Code: A Unified Framework for Evaluating Code Generation Across Tasks

Abstract
Advancements in LLMs have enhanced task automation in software engineering; however, current evaluations primarily focus on natural language tasks, overlooking code quality. Most benchmarks prioritize high-level reasoning over executable code and real-world performance, leaving gaps in understanding true capabilities and risks associated with these models in production. To address this issue, we propose MERA Code, a new addition to the MERA benchmark family, specifically focused on evaluating code for the latest code generation LLMs in Russian. This benchmark includes 11 evaluation tasks that span 8 programming languages. Our proposed evaluation methodology features a taxonomy that outlines the practical coding skills necessary for models to complete these tasks. The benchmark comprises an open-source codebase for users to conduct MERA assessments, a scoring system compatible with various programming environments, and a platform featuring a leaderboard and submission system. We evaluate open LLMs and frontier API models, analyzing their limitations in terms of practical coding tasks in non-English languages. We are publicly releasing MERA to guide future research, anticipate groundbreaking features in model development, and standardize evaluation procedures.
MERA Code: A Unified Framework for Evaluating Code Generation Across Tasks
Artem Chervyakov1, Alexander Kharitonov1, Pavel Zadorozhny1, Adamenko Pavel1, Rodion Levichev1, Dmitrii Vorobev1, Dmitrii Salikhov1, Aidar Valeev1,8, Alena Pestova3, Maria Dziuba2,3, Ilseyar Alimova7, Artem Zavgorodnev4, Aleksandr Medvedev4, Stanislav Moiseev4, Elena Bruches6, Daniil Grebenkin6, Roman Derunets6, Vikulov Vladimir5, Anton Emelyanov1, Vladimir V. Ivanov8, Dmitry Babayev1, Valentin Malykh3 Alena Fenogenova1 1 SberAI, 2 ITMO University, 3 MWS AI, 4 T-Technologies, 5 Rostelecom, 6 Siberian Neuronets, 7 Skoltech, 8 Innopolis University Correspondence: mera@a-ai.ru
1 Introduction
Recent advances in large language models (LLMs) have demonstrated significant potential for automating software engineering tasks, such as code generation and documentation. However, current evaluation methods predominantly emphasize natural language understanding, often neglecting critical factors like code quality, real-world applicability, and multilingual support. Benchmarks such as HumanEval-X Zheng et al. (2023b), MultiPL-E Cassano et al. (2022), and mxEval Athiwaratkun et al. (2023) address multilingualism through translated datasets but mainly focus on programming languages. They do not fully capture the interaction between natural language (such as requirements and comments) and code, which is vital for real-world development. This limitation restricts our understanding of LLMs’ practical capabilities, particularly in non-English contexts. Tasks such as writing localized documentation or interpreting vague requirements necessitate proficiency in both natural and programming languages, yet no comprehensive benchmark currently evaluates these cross-disciplinary skills. To fill this gap, we introduce MERA Code — a comprehensive evaluation system explicitly designed for the Russian language. MERA Code provides tools to assess LLMs in realistic, practical, multilingual software development scenarios.
Our key contributions are:
  • A reproducible evaluation methodology for LLMs in Russian;
  • A suite of 11 instruction-formatted tasks (code2text, text2code, code2code) across 8 programming languages (Python, Java, C#, JavaScript, Go, C, C++, and Scala);
  • An open evaluation platform with a scoring system, framework  [^1] The evaluation code https://github.com/MERA-Evaluation/MERA_CODE released under MIT License, and public leaderboard;
  • A set of performance analyses ranging from open-source general models to proprietary coding APIs.
MERA Code serves as a foundational resource for the research and industrial community, promoting collaboration to enhance task coverage and adapt to evolving LLM capabilities. By combining natural and programming language evaluation, we support more relevant assessments of LLMs in software engineering.
2 Related Work
Early benchmarks for evaluating LLM coding capabilities targeted simple, single-function problems with automated test-execution scoring, as exemplified by HumanEval Chen et al. (2021a) and MBPP Austin et al. (2021a), and by datasets drawn from online platforms such as APPS Hendrycks et al. (2021), CodeContests Li et al. (2022a), and their extension in TACO Li et al. (2023). As model context windows expanded, benchmarks introduced tasks of increasing scope: ClassEval Du et al. (2023) required correct implementations at the class level, while CoderEval Du et al. (2023) and RepoBench Liu et al. (2023b) assessed end-to-end project and repository proficiency. To better mirror developer workflows, SWE-Bench Jimenez et al. (2024) curated tasks from real GitHub issues, and BigCodeBench evaluated LLMs’ ability to orchestrate multiple function calls in practical programming scenarios. However, widespread dataset reuse has engendered significant data leakage, prompting the creation of dynamic evaluation suites such as LiveCodeBench Jain et al. (2024) and CodeElo Quan et al. (2025), which, despite mitigating test-set contamination, remain distant from the complexity of industrial software development.
Most existing benchmarks concentrate predominantly on code generation, overlooking the broader spectrum of development tasks such as code repair, execution, and test output prediction that are integral to real-world programming. Initiatives like CodeXGLUE Lu et al. (2021) span fourteen datasets covering ten tasks from defect detection to summarization and translation, and LongCodeArena Bogomolov et al. (2024) further incorporates self-repair and runtime evaluation. While multilingual programming benchmarks (e.g., HumanEval-X Zheng et al. (2023b), MultiPL-E Cassano et al. (2022), mxEval Athiwaratkun et al. (2023)) address various codebases, they generally neglect the multilingual natural-language elements of software engineering—requirements elicitation, documentation, commenting, and review. Consequently, no existing benchmark comprehensively measures LLM performance across the full software development lifecycle and in the diverse natural and programming languages encountered in practice.
3 MERA Code
3.1 Overview
MERA Code introduces a practical evaluation framework that expands the Family of MERA benchmarks [^2] https://mera.a-ai.ru/en/code Fenogenova et al. (2024) for assessing code-oriented and general-purpose state-of-the-art (SOTA) models. This framework serves as a foundational resource for both the research and industrial communities, fostering collaboration to improve task coverage and adapt to the evolving capabilities of LLMs. The release includes the proposed taxonomy of the model’s skills needed to solve practical coding tasks, along with a set of public and private test tasks in instruction-based format, an open evaluation codebase, and a web platform featuring an open automatic scoring system and leaderboard.
3.2 Methodology
3.2.1 Taxonomy
Several works proposed various taxonomies to systematize the field of software development and engineering. In Zhang et al. (2023), the taxonomy aligns with the software development pipeline and considers tasks that must be solved to deliver the product. The work Ruf et al. (2015) analyzes the field of software engineering from a perspective of programming skills required to solve specific tasks.
Our taxonomy is based on the latter approach since it allows us to decompose arbitrary tasks into a limited number of skills. Hence, the resulting taxonomy could be exhaustive yet straightforward.
A language model can be considered as a system consisting of inputs, internal state, and outputs. According to these parts, we derive four fundamental skills as a ground of our taxonomy: Perception (input), Reasoning and Knowledge (internal state), and Generation (output). All other skills form a tree, becoming increasingly niche with each new level. The taxonomy  [^3] We recognize that the current tasks in MERA do not encompass the complete taxonomy of skills. This serves as a guide for the community to understand what is covered and what needs to be added to future tasks in a systematic manner. presented in Figure 1 is based on Ruf et al. (2015), Zhang et al. (2023), and the authors’ domain expertise.
Figure 1: Taxonomy of MERA Code encompassing four foundational skills Perception / Knowledge / Reasoning / Generation utilized in the model to address certain tasks. Detailed explanation of each skill could be found in Appendix B.
3.2.2 Prompts selection
LLMs’ performance might significantly depend on a given prompt Shin et al. (2020), Gao et al. (2021), Fenogenova et al. (2024). To ensure an unbiased and robust assessment, each of the 11 tasks is paired with a set of distinct prompts. These prompts differ in syntactic framing, level of detail in the task description, and specification of the required output format. Prompts are assigned uniformly at random across all samples for a given task, with exactly one prompt applied per sample and the assignment is fixed. This strategy mitigates any inadvertent advantage or disadvantage conferred by a particular prompt style or model family affinity. Every prompt comprises a clear template for the expected output, thereby aligning model responses with the automated scoring system. The examples of prompts for each task could be found in Appendix A.
3.2.3 Evaluation procedure
All MERA Code tasks employ purely generative and instructive assessment where models continuously emit tokens until meeting predefined stopping conditions tuned to balance creativity, determinism, and safety. Each raw output undergoes task-specific post-processing to conform with metric requirements, including mandatory extraction of markdown-fenced code blocks from responses (with fallback to full outputs when parsing fails) before being scored with metrics. MERA Code utilizes the following metrics:
Pass@k Chen et al. (2021a) evaluates the functional correctness of the generated code.
Compile@k - evaluates the correctness of the compilation of the generated code.
chrF Popović (2015) is used in code-to-text tasks since its enhanced sensitivity to Russian morphological complexity.
BLEU Papineni et al. (2002a) measures the n-grams level similarity of the prediction and gold answer.
CodeBLEU (CBLEU) Ren et al. (2020a) combines regular BLEU with the measure of similarity of code syntax via abstract syntax trees (AST) and code semantics via data-flow.
Exact Match(EM) is the rate at which the predictions exactly match the true references.
Judge@k measures whether one of the top-k predictions matches with the gold answer via LLM-as-a-Judge Zheng et al. (2023a).
Total Score and Private Score are calculated as the mean value across all tasks and private tasks, respectively. For tasks that have multiple metrics, these metrics are also averaged.
3.3 Tasks
The MERA Code currently encompasses 11 code tasks. The statistics are presented in Table 1. We will briefly discuss the tasks listed below. Additional details and examples can be found in Appendix A.
CodeLinterEval
is a benchmark dataset designed to evaluate Python code generation and correction abilities based on linter errors, assessing models’ understanding of error messages, adherence to PEP 8 style, and preservation of code logic. It contains 110 samples with source code, linter feedback, instructions, and canonical solutions, using pass@k as an evaluation metric.
CodeCorrectness
is a benchmark for evaluating LLMs’ ability to predict whether unit tests compile and execute successfully or fail. The dataset consists of 1,361 samples written in Java, Python, and Go, which were collected from both human and LLM sources and subsequently selected based on automated filtering criteria. Evaluation is performed via EM against execution-verified ground truth.
RealCode and RealCodeJava
are benchmarks for evaluating the ability of code generation models to synthesize function bodies within real-world Python and Java projects, respectively. RealCode includes 802 tasks from 95 Python repositories created in 2024, while RealCodeJava comprises 298 tasks from 27 Java repositories created between 2024 and 2025. In both cases, tasks are constructed by identifying functions covered by tests, replacing their bodies with mocks, and retaining only those where tests fail on the mock but pass on the original. This ensures that the task requires generating semantically correct and test-sensitive logic. For each task, the model must generate the full function body based on the surrounding code context. Generated completions are inserted back into the project and evaluated automatically using the repotest [^4] https://github.com/MERA-Evaluation/repotest library and the project’s original test suite. The primary metric is pass@1.
JavaTestGen
is a benchmark for evaluating code generation models on the task of producing executable JUnit 5 unit tests for real-world Java classes. The benchmark consists of 227 tasks mined from open-source GitHub repositories. Each task provides the source code of a focal Java class, its name, and the expected test class name; the model must generate a complete and executable test file. Evaluation is fully automated: compile@1 measures the proportion of generations that compile successfully, while pass@1 quantifies those that pass all tests when executed with mvn test inside a provided Docker environment. We also use repotestfootnotemark: to run the tests and measure correctness.
StRuCom
Dziuba and Malykh (2025) is the first benchmark for evaluating LLMs’ ability to generate structured Russian-language code documentation. It comprises 153K examples across Python, Java, JavaScript, C#, and Go, sourced from human-authored GitHub repositories and synthetically generated examples. The $500$ samples were carefully selected from these examples as a test set for MERA Code. Evaluation employs the chrF metric due to its sensitivity to Russian morphological complexity.
YABLoCo
Valeev et al. (2025) is a benchmark that evaluates the performance of LLMs on large repositories. It focuses on C and C++ languages since other benchmarks do not fully cover them. YABLoCo consists of $208$ carefully selected functions from four large open-source repositories (bullet3, Redis, OpenSSL, LLVM). Each function is assigned a context category determined by the function’s dependencies. The primary purpose of the benchmark is to evaluate the quality of LLMs’ generations when various repository context is given alongside the prompt. The quality of the generated code is measured by the pass@1 and EM metrics.
RuCodeReviewer
The first benchmark for automated code review comment generation in Russian is designed to address the limitations of English-centric and often noisy datasets. The benchmark is built from a high-quality dataset of 689 merge-request diff-comment pairs sourced from an industry-oriented educational program, where professional developers review student code (Java, Python, Go, Scala). A rigorous two-stage curation process is employed, involving an LLM-assisted filter followed by verification from two independent human experts, to ensure that every comment is self-contained, actionable, and reproducible from the code diff alone. As an evaluation metrics, we use Judge@k, chrF, BLEU in a reference-based setup.
UnitTests
is a multilingual dataset for evaluating model abilities to generate unit tests for functions and methods in five programming languages: Python, Java, Go, JavaScript, and C#. It is built from GitHub repositories with open licenses. Each observation consists of a focal function/method, a unit test for it (test function/method), as well as the focal and test contexts collected from the repository for the focal and test functions, correspondingly. The goal of the model is to generate a unit test based on the provided focal function/method and its context, as well as the test function/method context. The evaluation metric is CBLEU.
ruHumanEval and ruCodeEval
are presented in the MERA benchmark [^5] https://mera.a-ai.ru/ru/text/tasks as public and private tests, respectively. In this version, we refine the prompt’s base, making it more precise for code completion tasks and eliminating ambiguous formulations that cause issues for models generating from scratch. We also change the generation processing algorithm to ensure better code blocks parsing and increase the maximum amount of new tokens to accommodate for models with higher reasoning abilities. The evaluation is conducted via pass@k metric.
Task name Language Metrics Size Prompts Skills
Private ruCodeEval Python pass@k 164 10 Instruction Following, Code Perception, Completion, Algorithms & Data Structures
RuCodeReviewer Java, Scala, Go, Python Judge@k, BLEU, chrF 689 10 Instruction Following, Code Perception, Review, Simulation, Explanation, Programming Patters, Style Guides
CodeLinterEval Python pass@k 110 10 Instruction Following, Code Perception, Style Guides, Review, Editing
Public ruHumanEval Python pass@k 164 10 Instruction Following, Code Perception, Completion
StRuCom Python, Java, Go, C#, JavaScript chrF 500 10 Instruction Following, Code Perception, Simulation, Documentation
UnitTests Python, Java, Go, C#, JavaScript CodeBLEU 2500 20 Instruction Following, Code Perception, Synthesis, Testing, Long Context Comprehension
CodeCorrectness Python, Java, Go EM 1361 11 Instruction Following, Code Perception, Simulation, Error Classification
RealCode Python pass@k 802 10 Instruction Following, Code Perception, Completion
RealCodeJava Java pass@k 298 10 Instruction Following, Code Perception, Completion
JavaTestGen Java pass@k, compile@k 227 10 Instruction Following, Code Perception, Completion, Testing
YABLoCo C, C++ pass@k, EM 208 11 Instruction Following, Code Perception, Completion, Long Context Comprehension
Table 1: The MERA Code tasks outline. Size is the number of test samples. Prompts is the number of unique prompts used for each task. Skills lists skills from taxonomy which are necessary for particular task.
3.4 System Demo
The automatic scoring system is accessible on the benchmark platform  [^6] https://mera.a-ai.ru/en/code. The submission process involves the following steps. First, users must clone the MERA Code benchmark repository  [^7] The MERA Code is based on the fork of the LM-evaluation harness Gao et al. (2024). and create submission files using a shell script and the provided customized evaluation code. Second, they must register on the benchmark platform and upload their submission files through the interface in their account for automatic evaluation. The scoring process may take up to three hours. Once the evaluation is complete, the results are displayed in the user’s account and remain private unless the user opts to publish them using the “Publish” function. In this case, the submission undergoes expert verification for reproducibility, which includes checking the log files automatically generated by the evaluation script and reviewing the provided submission information. Once approved, the model’s score is publicly visible on the leaderboard, while the specific outputs remain private. The user process for the benchmark is illustrated in Figure 2 of the Appendix.
4 Evaluation
The MERA Code evaluation procedure is designed to strike a balance between accessibility and rigorous, standardized assessment of code generation capabilities. Users begin by cloning the benchmark’s GitHub repository, built on the lm-evaluation-harness framework  [^8] https://github.com/EleutherAI/lm-evaluation-harness with officially integrated customizations, and execute a dedicated script with fixed hyperparameters to launch lm-eval for a specified model and task set. Two repository branches are available: one enables local scoring of public tasks with reference outputs, while the other excludes complex dependencies to facilitate submission to the automated scoring system. While users may choose either workflow, local scoring does not guarantee metric correctness for tasks dependent on runtime checks or hidden test cases. Generated outputs are saved in JSON files and submitted as a ZIP archive via the MERA Code website, along with additional information to ensure reproducibility. The automated system then processes submissions, performing runtime evaluations that may take hours, and delivers detailed metrics to the user’s account page.
4.1 Baselines
We evaluate both code-oriented and general-purpose open models of various sizes: Qwen2.5 Coder (32B-Instruct), Qwen2.5 (72B-Instruct), Deepseek Coder V2 (236B-Instruct), ByteDance Seed-Coder (8B-Instruct), Mixtral (8x22B-Instruct), Yi Coder (9B-Chat), and Vikhr-YandexGPT-5-Lite (8B-it). Additionally, we assess proprietary models such as OpenAI GPT-4  [^9] https://openai.com/api/, Gemini 2.5  [^10] https://deepmind.google/models/gemini/, GigaChat 2 Max  [^11] https://giga.chat/, GigaCode  [^12] https://gigacode.ru/.
4.2 Results
Model Total Score Private Score
YABLoCo
RealCode
RealCodeJava
ruCodeEval
ruHumanEval
JavaTestGen
CodeLinterEval
ruCodeReviewer
UnitTests
StRuCom
CodeCorrectness
pass@1 Judge@1 CBLEU chrF EM
GPT-4.1 0.377 0.382 0.144 0.418 0.416 0.443 0.45 0.344 0.555 0.096 0.162 0.297 0.66
GPT-4o 0.377 0.381 0.149 0.324 0.399 0.529 0.537 0.37 0.479 0.078 0.153 0.275 0.666
Gemini 2.5 flash 0.356 0.427 0.12 0.388 0.386 0.61 0.604 0.211 0.496 0.064 0.174 0.217 0.404
DeepSeek-Coder-V2-Inst 0.347 0.36 0.149 0.363 0.383 0.433 0.392 0.269 0.494 0.06 0.153 0.2 0.714
GigaChat 2 Max 0.346 0.365 0.106 0.332 0.342 0.537 0.53 0.137 0.425 0.062 0.175 0.294 0.68
Qwen2.5-Coder-32B-Inst 0.296 0.306 0.111 0.323 0.383 0.311 0.289 0.22 0.466 0.065 0.168 0.213 0.519
GigaCode 1.4 0.289 0.166 0.135 0.322 0.352 0.357 0.305 0.313 0.027 0.064 0.189 0.276 0.676
Qwen2.5-72B-Inst 0.285 0.254 0.144 0.362 0.349 0.174 0.157 0.189 0.481 0.048 0.128 0.252 0.702
Seed-Coder-8B-Inst 0.268 0.345 0.106 0.106 0.305 0.317 0.21 0.264 0.655 0.035 0.128 0.237 0.403
Yi-Coder-9B-Chat 0.203 0.181 0.135 0.067 0.252 0.35 0.173 0.229 0.145 0.016 0.138 0.192 0.364
Mixtral-8x22B-Inst 0.179 0.028 0.106 0.18 0.366 0.0 0.0 0.229 0.027 0.016 0.159 0.152 0.597
Vikhr-YandexGPT-5-Lite-8B 0.168 0.187 0.091 0.032 0.201 0.035 0.024 0.123 0.407 0.022 0.106 0.138 0.464
Table 2: MERA Code benchmark results. The private tasks are CodeLinterEval, ruCodeEval and ruCodeReviewer.
Table 2 shows the results of the baselines evaluation. Gemini 2.5 Flash (0.356), GPT-4o (0.377) and GPT-4.1 (0.377) demonstrate superior overall performance across eleven coding tasks, particularly excelling in multi-language documentation (StRuCom), Python/Java completions (RealCode/RealCodeJava/ruHumanEval/ruCodeEval), and unit-test generation. They are closely trailed by DeepSeek-Coder-V2 (0.347) and GigaChat-2-Max (0.346), with the former showing exceptional capability in predicting code compilation success (CodeCorrectness: 0.714). Vikhr-YandexGPT-5-Lite lags significantly (0.168), showing near-zero performance in algorithmic and code completion tasks. All models exhibit pronounced weaknesses in generating unit-tests on Python (UnitTests) and automated comments generation (ruCodeReviewer), where maximum scores remain below 0.1 despite task apparent simplicity.
Specialized models show targeted strengths: DeepSeek-Coder-V2 excels in compilation prediction (0.714 vs GPT-4o’s 0.666), while Seed-Coder-8B dominates error correction (0.65 for CodeLinterEval - about 35% better than the average result). Mixtral-8x22B outperforms GPT-4o on multilingual unit-tests generation (+4%). Qwen2.5-72B demonstrates worse results compared to Qwen2.5-Coder-32B (-4-20% on half of the tasks) differing the most on comments generation (-25%), predicting code executability (+28%), code generation on C/C++ (+23%) and algorithmic tasks (-45%). GigaChat-2-Max shows strength in complex Python completions (RealCode: 0.332, ruCodeEval: 0.537), and GPT-4.1 maintains consistent documentation quality (StRuCom: 0.297). GigaCode 1.4 achieves the best score in the generation of multilingual unit-tests (0.188). However, inconsistent JUnit5 test generation highlights critical gaps. These results affirm that while versatile architectures like GPT-4 excel broadly, targeted fine-tuning yields niche advantages—though reasoning and debugging capabilities remain key challenges across all models. Extended results could be found in Appendix C.
5 Conclusion
In this work, we present the MERA Code, a comprehensive evaluation tool designed for coding tasks and programming languages using Russian prompts across eight different languages. MERA provides an open-source evaluation toolkit, a structured scoring system, and a public leaderboard, enabling standardized assessments of coding skills while addressing limitations in non-English contexts. This framework enhances our understanding of model capabilities and helps identify strengths and weaknesses based on community tests. We are releasing MERA to the research community to encourage innovation, establish standards, and promote reproducible evaluation practices.
Limitations
While the proposed benchmark advances the evaluation of LLMs’ coding abilities, several limitations remain.
Limited Representativeness
Despite efforts to ensure coverage, our datasets cannot comprehensively represent the entire landscape of programming problems, especially as real-world coding tasks are highly diverse and domain-specific. As a result, while our benchmark captures a broad range of difficulty levels and programming domains, certain edge cases and emerging languages or paradigms may not be well represented. Technical challenges related to non-Python languages complicate our ability to cover this variety.
Code Quality Assessment
We employ various types of evaluations across the datasets, including n-gram matching, LM-as-Judge, and pass@k. While many datasets utilize the LM-as-Judge technique for evaluation, this method does not always adequately assess deeper aspects of code quality, such as readability, efficiency, maintainability, security, and adherence to best practices. Additionally, significant challenges arise from the parsing of generated answers and the models’ instruction-following capabilities, which can affect the accuracy of the results. These qualitative factors often require more nuanced human judgment, which cannot be easily scaled for everyday benchmarking.
Testing Conditions
The benchmark currently assumes that coding prompts are well-specified and unambiguous, which is often not the case in practice. The ability of large language models (LLMs) to handle unspecified, noisy, or ambiguous real-world requirements remains a challenge. Additionally, our evaluation infrastructure assumes that the generated code can be executed in isolated environments, which may not accurately reflect the complexity and constraints of production development settings.
There are instances where Docker containers cannot be built due to issues arising from the network setups of several datasets (such as RealCode, RealCodeJava, and JavaTestGen) that are designed to reproduce the conditions found in their respective repositories. These containers may not always compile successfully. While we strive for accuracy in our deployment, we cannot guarantee that it will always score everything perfectly. However, any discrepancies are typically within 4%, which accounts for about 0.2% of the total, and should not significantly impact the overall ranking.
Data Contamination
While we have implemented measures to minimize data contamination (e.g., by removing public problems found in known training sets), there remains the possibility that LLMs — particularly when trained on vast internet corpora — may have encountered similar or even identical problems during training, potentially inflating performance results. Additionally, some datasets were created from repositories that might contribute to the risk of data contamination.
Scoring optimization
The scoring system in the benchmark takes time to produce results, often requiring up to 2 hours due to the complexity of running environments and libraries for various programming languages. We plan to address code optimization and testing environments in the future.
Technical Constraints
As with any rapidly evolving technology, improvements in LLM architectures and training data may quickly impact the relevance of our benchmark. Therefore, regular updates and community engagement will be necessary to maintain its utility as a robust evaluation tool.
Ethical Statement
As with any research that advances the development and evaluation of LLMs, introducing a new coding benchmark necessitates careful consideration of ethical implications.
Firstly, the benchmark has been constructed mindful of data privacy and intellectual property rights and released under the MIT licence. All programming tasks and datasets utilized are either original or derived from public domain resources, and are used with proper attribution and permissions. We have taken precautions to ensure that example problems do not contain proprietary or sensitive information and that datasets do not inadvertently leak private user data.
Secondly, the public release of such a benchmark may facilitate broader research into improving LLMs for software development. While this has clear benefits in promoting open scientific progress, it may also be misused, for example, to enhance automated systems capable of generating malicious code or exploiting software vulnerabilities. We strongly discourage the use of this benchmark, or any resulting models, for unethical or harmful purposes and urge the community to adhere to responsible usage guidelines.
Thirdly, although LLMs have the potential to democratize programming by lowering barriers to software development, concerns exist regarding their impact on the workforce, code quality, and reliance on automated tools. Our benchmark is intended strictly for research and evaluation purposes. We encourage users to complement model assessment with human oversight and to remain vigilant for biases, errors, or unsafe behavior that may emerge from automatic code generation.
Lastly, by releasing the benchmark and associated evaluation code, we commit to transparency and reproducibility in our research. We invite feedback from the broader community, particularly regarding unintended biases, fairness across programming languages, and the benchmark’s accessibility for a global and diverse audience.
AI-assistants Help We improve and proofread the text of this article using Writefull assistant integrated into Overleaf (Writefull’s/Open AI GPT models), Grammarly [^13] https://app.grammarly.com/ to correct grammatical, spelling, and style errors and paraphrase sentences.We want to clarify that these tools are used exclusively to improve the quality of English writing, in full accordance with ACL policies regarding the responsible use of AI writing assistance. However, some parts of our publication may potentially be identified as AI-generated, AI-edited, or a combination of human and AI contributions.
Acknowledgments
We gratefully acknowledge the AI Alliance Russia [^14] https://a-ai.ru/?lang=en for organizing and supporting our work, offering legal guidance, and providing computational resources essential for the development and maintenance of the MERA Code benchmark.
We also extend our sincere thanks to all our industrial and academic partners for their valuable contributions to the MERA Code — through their shared expertise and datasets, they have made this benchmark possible.
We would like to express our deep appreciation to Ekaterina Morgunova, Egor Nizamov, Ivan Bondarenko, Georgy Mkrtchyan, Ivan Kharkevich, Nikolay Bushkov, Irina Shakhova, Denis Kokosinsky, Kirill Pikhtovnikov, Anton Bykov, and Oleg Sedukhin for their contributions and generous support throughout this work.
In the following list, we mention all authors of MERA Code with a description of their contribution:
  • Codebase techlead & Contributor of ruCodeEval and ruHumanEval & Baseline evaluation & Deployment: Artem Chervyakov,
  • Methodological Review & Taxonomy of MERA Code & Code review & Baseline evaluation & Deployment: Alexander Kharitonov,
  • Contributors of RealCode, RealCodeJava, JavaTestGen: Pavel Zadorozhny, Adamenko Pavel, Rodion Levichev, Dmitrii Vorobev, Dmitrii Salikhov ,
  • Contributor of YABLoCo: Aidar Valeev,
  • Contributor of UnitTests: Alena Pestova,
  • Contributor of StRuCom: Maria Dziuba,
  • Contributors of ruCodeReviewer: Ilseyar Alimova, Artem Zavgorodnev, Aleksandr Medvedev, Stanislav Moiseev,
  • Contributors of CodeCorrectness: Elena Bruches, Daniil Grebenkin, Roman Derunets,
  • Research advisor & Contributor of CodeLinterEval: Vikulov Vladimir,
  • Baseline evaluation & Deployment: Anton Emelyanov,
  • Research advisor & Contributor of RealCode, RealCodeJava, JavaTestGen: Dmitrii Babaev,
  • Research advisor & Contributor of YABLoCo: Vladimir V. Ivanov,
  • Research advisor & Contributor of StRuCom and UnitTests: Valentin Malykh,
  • Administration & Ideology & Contributor of ruCodeEval and ruHumanEval & Senior research advisor: Alena Fenogenova
Appendix A Datasets examples
ruHumanEval
ruHumanEval is the Russian counterpart of the HumanEval dataset Chen et al. (2021a), which assesses models’ abilities to generate solutions for straightforward programming problems in Python. The dataset contains the translated into Russian and manually verified tasks of the original dataset  [^15] https://huggingface.co/datasets/openai_humaneval, including the test cases, which were taken from Liu et al. (2023a) (10 test cases per task).
ruCodeEval
ruCodeEval is created from scratch by assembling various programming tasks of the same difficulty level as the ruHumanEval test part and manually writing the test cases and documentation strings. All tasks were verified to ensure that no repetition (with ruHumanEval) occurred in the test samples.
These tasks evaluate the functional correctness of code generation by providing input information, including a textual function description (docstring) and examples of expected results for different test cases.
An example of ruCodeEval and ruHumanEval tasks:
  • instruction: The input represents a function with a description in the form of a docstring. Given the input function, you need to implement it based on the template: ‘‘{function}’’.
  • function:
    def gcd(a: int, b: int) -> int:
    """Returns the greatest common divisor of two integers a and b.
    Examples:
    gcd(3, 5)
    1
    gcd(25, 15)
    5"""
  • tests: "[{’a’: 3, ’b’: 7}, {’a’: 10, ’b’: 15}, {’a’: 49, ’b’: 14}, {’a’: 144, ’b’: 60}]"
  • outputs (golden answer): [1, 5, 7, 12]
JavaTestGen
To construct JavaTestGen, we first collected 5,000 of the most starred and recently updated Java repositories on GitHub. Repositories without an open-source license, larger than 100MB, or those that failed to compile and pass their tests using Maven were excluded. From the remaining pool, we selected repositories with at least two passing test cases and fully green test suites, yielding 500 high-quality Java projects.
Using static analysis, we extracted focal classes and their corresponding test classes, retaining only self-contained test modules (i.e., those depending solely on the focal class’s module). We further filtered class pairs with the following criteria: (i) source file length between 1,000 and 5,000 characters; (ii) not an enum or interface; (iii) at least two public methods in the source class; and (iv) test class with at least two assertions. This resulted in a curated set of 227 focal–test pairs.
At inference time, the model is given only the focal class and must generate the corresponding JUnit 5 test class. For evaluation, we replace the original test file with the generated one and run mvn clean test in a controlled Dockerized environment. We report two metrics: compile@1, the proportion of generated tests that compile, and pass@1, the proportion that pass all assertions.
An example of JavaTestGen task:
  Context code:
package com.github.quiram.course;
import java.util.List;
import static java.lang.String.join;
import static java.util.Arrays.asList;
public class ReverseCommand extends
Command {
    @Override
    protected boolean
    safelySupports(String input) {
  Prompt Template (translated)
*This prompt was used during data generation and translated from Russian:
You are given the implementation of a class {class}
{context code}
Your task is to write a unit test class {test_class} for the class above.
Use Junit5. Ensure full coverage of all test scenarios even if the code doesn’t contain corresponding branches. Write tests for happy path, edge cases, illegal arguments and other cases. Ensure only one assert in each test method. Test method names should be meaningful. Add necessary imports and annotations.
RealCode
Recent work has demonstrated that widely used surface-level metrics for code generation, such as BLEU Papineni et al. (2002b) and CodeBLEU Ren et al. (2020a), correlate poorly with functional correctness and real-world utility Chen et al. (2021b); Dong et al. (2025). These metrics tend to reward superficial similarity to reference implementations, often failing to capture whether the generated code satisfies the intended behavioral requirements Dong et al. (2025). Consequently, leading benchmarks such as HumanEval Chen et al. (2021b), MBPP Austin et al. (2021b), and AlphaCode (ClassEval) Li et al. (2022b) rely on execution-based evaluation, using pass@k (a task is considered solved if any of $k$ samples passes all tests) as the primary correctness metric. However, these benchmarks are typically limited to synthetic or isolated tasks, reducing their ability to reflect practical code completion and maintenance scenarios Zhuo et al. (2025).
To address these limitations, we introduce RealCode, a benchmark designed to evaluate code generation models on realistic code completion tasks within authentic, actively maintained Python repositories. We curated projects from public GitHub repositories created in 2024, filtering for open-source licenses, a minimum of three GitHub stars and one fork, and repository size under 30MB to ensure real-world relevance and manageable computational overhead. Additionally, we excluded repositories where function-level docstrings were consistently missing or empty, as documentation is often essential for guiding meaningful code generation. Each repository was automatically cloned, built, and tested using a rule-based pipeline in a Dockerized environment, retaining only those that successfully built and passed their test suites.
We then identified all functions within each project that were covered by at least one test. For these functions, we replaced the body with a mock implementation (e.g., a pass statement or a default return value) and reran the project’s test suite. We retained only those functions for which the tests failed with the mock but passed with the original implementation, ensuring that solving the task requires synthesizing semantically meaningful and test-sensitive code.
Each task is defined by the left-side context (i.e., the beginning of the file and the function signature up to the body), with the target being the original function body. The model is expected to generate only the function body, which is inserted back into the codebase during evaluation.
The generated completion is then integrated into the original repository and evaluated using pytest with the project’s existing test suite. The primary evaluation metric is pass@1, indicating whether the generated code passes all relevant tests. All steps, from project preparation to test execution, are fully automated and reproducible via our open-source repotest library.
Prompting Considerations. One of the challenges in code generation benchmarks is enforcing consistent formatting in model outputs. Generated completions must be syntactically valid and correctly indented to be parsable and executable in the original project. This is especially important in languages like Python, where whitespace is semantically meaningful. Our prompt format ensures that models produce completions that integrate cleanly into the original context. The second major challenge is semantic correctness: even if the format is valid, the model must synthesize correct logic that satisfies the tests based on limited surrounding context.
An example of RealCode task:
  Left Context:
from dataclasses import dataclass
from typing import Self
from pysphinx.const import SECURITY_PARAMETER
from pysphinx.crypto import compute_hmac_sha256
@dataclass
class IntegrityHmac:
    This class represents a HMAC-SHA256
    that can be used
    for integrity authentication.
    value: bytes
    SIZE: int = SECURITY_PARAMETER
    def __init__(self, value: bytes):
        """Override the default constructor
        to check the size of value"""
  Ground Truth Body:
        if len(value) != self.SIZE:
            raise ValueError("invalid length of HMAC",
                            len(value))
        self.value = value
  Prompt Template (translated)
*This prompt was used during data generation and translated from Russian:
Use the following code: {left_context}
Now — only implement the body of a single function. Wrap your answer in a block:
‘‘‘python
<code>
Indentation must be correct. Do not include the function signature. Do not write other functions. Your answer will be inserted into the function and tested.
RealCodeJava
The benchmark aims to evaluate the code generation model’s ability to produce compilable and functional Java code in the context of genuine open-source Java repositories.
Tasks were gathered from public GitHub repositories created in 2024–2025, which were thoroughly filtered to include only repositories with open-source licenses and those that met popularity and utility criteria (minimum 3 stars on GitHub). We then selected projects that used the Maven build system and had unit tests within. We left only those that could be successfully built and that had passed the tests.
Each task consists of a function, which is covered by at least one test. We verify the test’s capability to detect non-working code by replacing the function body with an auto-generated mock implementation. Specifically in Java, our mocks return null values, empty collections, or some predefined constants depending on the function’s signature. Mocks are expected to pass the compilation stage but fail during the test execution phase. We only keep samples that meet this criterion. Finally, all tasks are automatically evaluated for perceived target function complexity and ranked accordingly. We approximate the perceived complexity of a function by the presence of certain patterns in the code, such as cross-file API calls, usage of complex lambda expressions, or safeguard checks.
For each task, inputs include left-side context (the beginning of the file and the function signature with opening curly brace "{"). From there, the model is expected to generate the body of the function, namely the lines of code up to the last closing curly brace "}", including it. The resulting answer is then inserted back into the codebase and verified.
Generated completions are evaluated using the original project’s existing test suite and our open-source repotest library. We use pass@1 as the main metric to measure the model’s ability to provide a functional solution to the given task.
An example of RealCodeJava task:
  Left Context:
package com.comp301.a06image;
import java.awt.Color;
 * The SolidColorImage class is
 * an Image implementation
 * that represents an image with a solid color.
 * The image is defined by
 * a width, a height, and a color.
 * The color of a pixel can be retrieved by
 * calling the getPixelColor method.
public class SolidColorImage implements Image {
  private int width;
  private int height;
  private Color color;
    * Returns the color of the pixel at the specified
   * coordinates.
   * @param x The x coordinate of the pixel
   * @param y The y coordinate of the pixel
   * @return The color of the pixel
  @Override
  public Color getPixelColor(int x, int y) {
  Ground Truth Body:
    if (x < 0) {
        throw new IllegalArgumentException(
                "Pixel x-coordinate must be non-
                negative.");
    if (y < 0) {
        throw new IllegalArgumentException(
                "Pixel y-coordinate must be non-
                negative.");
    if (x >= this.width) {
        throw new IllegalArgumentException(
                "Pixel x-coordinate must be less than
             the image width.");
    if (y >= this.height) {
        throw new IllegalArgumentException(
                "Pixel y-coordinate must be less than
             the image height.");
    // Return the solid color of the image
    return this.color;
  Prompt Template (translated)
*This prompt was used during data generation and translated from Russian:
Use the following code: {left_context}
Write the contents of the last function after the header with arguments. Do not invent new functions and classes, but you can use existing ones in the code. The answer consists of one function. Place the answer in the block:
‘‘‘java
function body
Apply the indents and formatting as in the example.
RuCodeReviewer
The first benchmark for automated code review comment generation in Russian, designed to address the limitations of existing English-centric and often noisy datasets. The creation of this dataset is motivated by the need for a high-quality, reproducible evaluation framework for models’ abilities in the code review task.
The data for RuCodeReviewer originates from Backend Academy, an industry-oriented educational program where student-submitted code across four languages (Java, Python, Go, and Scala) is manually reviewed by professional software developers. Utilizing the GitHub API, 1,300 merge requests containing 2,859 review comments were collected. A rigorous two-stage pipeline was used to clean the data. First, an LLM-assisted filtering stage, based on GPT-4o, identifies comments that are self-contained and reproducible solely from the provided code diff. A manually validated subsample of 1,300 examples marked as "non-reproducible" at this stage contained no misclassified cases, demonstrating the high quality of the proposed approach. Second, two independent human annotators, both experienced software developers, verify each "reproducible" candidate, ensuring that comments are actionable and fully understandable without external context. This human validation step results in a substantial inter-annotator agreement, with a Cohen’s kappa coefficient of 0.78. The final dataset is consists of 689 merge-request diff-comment pairs prioritizing quality over size.
To overcome the “one-to-many” challenge in evaluation of the generation tasks, where multiple valid outputs exist for a single input, a novel evaluation methodology centered on an LLM-as-a-Judge was proposed. This framework leverages a large language model to assess the semantic equivalence between model-generated and human reference answers. Performance is reported using a pass@k metric, which credits a model if at least one of its $k$ generated candidates is deemed valid. The Qwen2.5-Coder-32B  [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-32B model was used as the judge model, so its performance measurements on this benchmark may be biased. To verify the accuracy of the LLM-as-a-judge metric, comments were generated using GPT-4o for 45 examples from the dataset. These comments were annotated by human annotators for semantic equivalence with the original comments provided by real reviewers. Based on these annotated examples, the prompt for the Qwen2.5-Coder-32B judge model was refined until 100% agreement with human annotators was achieved on these 45 examples. Subsequently, the model’s agreement with annotators was further evaluated on an additional 100 comments generated in the same way using GPT-4o, achieving a Spearman correlation of 0.831 with human annotators and confirming its reliability for this task.
An example of RuCodeReviewer task:
Enum values
  • Diff Block (Original):
    +CategoryType = tuple[tuple[str], tuple[str]]
    +class LetterStatus(Enum):
    +   NO_USED: int = 0
    +   USED_IN_WORD: int = 1
    +   USED_NOT_IN_WORD: int = 2
  • Gold comment (Translated): NO_USED = enum.auto()
Go benchmark loop
  • Diff Block (Original):
    +func BenchmarkSummation(b testing.B) {
    +  sizes := []int{config.HundredThousands, ...}
    +  for _, size := range sizes {
    +    data := make([]int, size)
    +    b.Run(fmt.Sprintf("chan_size_+    b.Run(fmt.Sprintf("mu_size_+    b.Run(fmt.Sprintf("wg_size_+  }
  • Gold comment (Translated): Add a for range b.N loop — for very fast functions, a single run might be insufficient.
Java null checks
  • Diff Block (Original):
    +import static backend.academy.Values.RANDOM_COLOR;
    +import static backend.academy.Values.VARIATIONS;
    +@Getter
    +public class Transformations {
    +  private Color lastColor;
    +  private final Coefficients[] coefficients;
    +  private final double[] weights;
    +  private final Color[] colors;
    +  private final PointFunction[] variations;
  • Gold comment (Translated):
    - Missing null checks for arrays
    - variations is not validated
Python nesting
  • Diff Block (Original):
    +    self._bytes_sum = 0
    +    self._filter_name = filter_name
    +  def parse(self) -> LogReport:
    +    Method to begin parsing data
    +    files = glob.glob(self._path)
  • Gold comment (Translated):
    Excessive nesting. Split checks into separate functions and simplify the block;
    currently it’s a loop → try/except → context manager → another loop with branching — it’s too much.
UnitTests
UnitTests is a multilingual dataset for evaluating models on generating unit tests in Python, Java, Go, JavaScript, and C#. Each example includes a function, its context, and the corresponding test with its context.
The data is sourced from open-licensed GitHub code. First, a list of repositories for each language was obtained. We chose the repositories with permissive licenses only and with the number of stars more than 10. We also filtered out fork repositories. The list of licenses used in the dataset: MIT License, Apache License 2.0, The Unlicense, Mozilla Public License 2.0, BSD 2-Clause "Simplified" License, BSD 3-Clause "New" or "Revised" License, EPL 1.0 license, EPL 2.0 license, MPL 2.0 License, Unlicense License, 0BSD license.
Then, the repositories were downloaded and parsed using syntax parsers. For all languages (except Python) the tree-sitter was used for code parsing, specifically, for searching and parsing functions/methods and classes, identifying calls, etc. For Python, we use the built-in ast library.
Next, methods and functions, along with their unit tests, were mapped using a method adopted from the paper Tufano et al. (2021). For Java, the mapping procedure was identical to the method by Tufano et al. (2021). For Python, Go, and C#, we develop similar mapping methods based on functions names and unique method invocation. For Javascript, the tests were mapped to functions by the last local method/function invocation in the test because test functions do not have identifiers when declared in it() and test().
When building the dataset, the same filtering rules for all languages were used: (i) Empty tests are removed. (ii) No more than 200 method-test pairs were collected from one repository. If there were more pairs, they were sampled randomly. (iii) The test case should be less than 5000 characters. This limit is set to remove overly long tests from the data. (iv) Maximum input length (focal function with context) should be less than 70000 characters. (v) Maximum number of assertions (the word "assert" in the test case) is 20. (vi) For Python and Java, there was additional filtering for tests with syntax errors (using ast and javalang libraries correspondingly). (vii) The training data was filtered for duplicates of test cases both within a set, and possible overlaps with the validation and test data were removed.
For the benchmark, the data was sampled for each language so as to cover different cases as evenly as possible in terms of the length of the focal function, the test function, and the entire context; and in terms of the type of both functions (function/method). The benchmark comprises a total of 2,500 samples, with 500 samples for each programming language. For evaluation, we use the CodeBLEU Ren et al. (2020b) metric to compare the original test from the repository and the generated test.
  • instruction: Write a test for the following go code from the file ’lists/mergesorted.go’.
    You need to write the test function on go. The test will be placed to the file ’lists/mergesorted_test.go’.
    You can use the following entities imported or declared in the test file:
    package lists
    import (
    "math/rand"
    "reflect"
    "sort"
    "testing"
    Pay attention to the following code when writing the test:
    #lists/mergesorted.go
    package lists
    #focal function/method here
    Code for testing:
    func MergeSorted(l, f *List) (m *List, ok bool)
    m = new(List)
    for l.Len() > 0 || f.Len() > 0
    vl, nl, okl := PopInt(l)
    vf, nf, okf := PopInt(f)
    if !okl || !okf
    return m, false
    ll, n := l, nl // The assumption is: vl <= vf.
    switch
    case l.Len() == 0:
    ll, n = f, nf
    case f.Len() == 0:
    ll, n = l, nl
    case vl > vf:
    ll, n = f, nf
    m.Insert(ll.Remove(n))
    return m, true
    Write only the test function without any explanations or comments.
    Your answer should be formatted using markdown as follows:
    ```go
    <your code>
CodeLinterEval
CodeLinterEval is a dataset for evaluating model abilities for generating and correcting code based on linter errors in the Python language. The benchmark evaluates understanding linter errors (the ability to correctly interpret messages like E111, E231, etc.), correct code refactoring (the ability to make corrections while preserving the logic of the program), following the code style (PEP 8) that includes correct indents, spaces, formatting, and contextual understanding of the code - the model should not break the logic when fixing the style. The benchmark contains 110 tasks: source code with errors, feedback – a list of errors with description from the linter, instruction – explicit instruction to correct the code based on feedback, and reference canonical code.
Explicit indication of errors allows us to evaluate the model’s ability to correct code according to the linter, rather than "guess" errors. The canonical solution provides a clear ground truth for evaluation. The instruction explicitly specifies the task so that the model does not deviate from the goal.
If the model corrects the code according to the feedback and the linter does not detect errors after checking the generation results, then the result is correct. If errors persist or new ones appear, the model does not solve the problem. The metric is pass@k, determined based on the success of the linter check relative to the total size of the dataset.
An example of CodeLinterEval task:
  • instruction: Rewrite the code based on the errors received from the linter. Errors indicate critical weaknesses: potential bugs, security vulnerabilities and violations of clean code principles. Fix ALL these errors without exceptions, keep the original logic of the program, strictly adhere to PEP-8 for Python, do not add comments and explanations. Errors and warning from the linter: “{feedback}” Code: “{code}”. Provide the response in the format corresponding to the template for the response:
    ``` python
    <code>
  • code:
    def first_repeated_char(str1):
      for index,c in enumerate(str1):
        if str1[:index+1].count(c) > 1:
          return c
  • feedback: E111: indentation is not a multiple of 4 in 2 line. E231: missing whitespace after ’,’ in 2 line. E111: indentation is not a multiple of 4 in 4 line. W292: no newline at end of file in 4 line
CodeCorrectness
CodeCorrectness is a novel benchmark designed to evaluate LLMs’ ability to identify whether a unit test is correct (compilable and executable) or incorrect (failing compilation or execution). Although advanced tools are being developed to assess generated code correctness Bui et al. (2025), dedicated benchmarks for evaluating LLMs’ capabilities in code assessment remain an underexplored research area.
The dataset comprises 1,361 samples across Java, Python, and Go, constructed through a multi-stage process. Focal file (the code under test) – test file pairs were first automatically collected from GitHub repositories using filters for permissive licenses (e.g., MIT, Apache-2.0), repository popularity (stars), and recency (last commit date). Exclusion criteria removed pairs with minimal test or focal code lines, additional imports from the project, imports from specific libraries, or file operations in test cases. The resulting samples include both 466 original human-written tests that passed all filters, compiled and executed successfully, alongside 895 LLM-generated tests. A subset of these LLM-generated tests contains errors, failing during compilation or execution; samples exhibiting only syntax errors were filtered out. This process ensured sufficient context for the task and provided real-world complexity.
We employ exact match as the evaluation metric for this task. Each sample presents a focal file code and a corresponding test file code containing test cases. Models are tasked with predicting a binary label: compilation and execution success or failure. Ground truth labels were rigorously verified using actual execution within isolated environments, confirming that a specific failure reason occurred for failing tests.
  • instruction:
    Below are the focal and test files. Please check if the test will run correctly.
    Focal file:
    ```{lang}
    {focal}
    Test file:
    ```{lang}
    {test}
    Your answer: if the test runs without errors, write <<success>> otherwise— <<failed>>.
  • language: Go
  • focal file (snippet):
    type Queue[V core.Value] struct {
        items []core.Node[V]
  • test file (snippet):
    func TestQueue_WithLargeNumberOfItems
                        (t *testing.T) {
        q := NewQueue[int]()
        assert.True(t, q.IsEmpty())
  • outputs: success
StRuCom
StRuCom Dziuba and Malykh (2025) addresses a critical challenge in AI-assisted code documentation by introducing the first large-scale dataset (153K examples) of structured Russian comments for Python, Java, JavaScript, C#, and Go. This resource enables robust evaluation of documentation generation models for Russian-speaking developers, filling the void of standardized benchmarks for non-English code explanations.
To construct StRuCom, the authors developed a hybrid sourcing methodology combining human-authored and AI-generated content. Human-written comments were extracted from over $150$K Russian GitHub repositories, identified through repository metadata analysis and license filtering. To ensure representation across all target languages, the corpus was supplemented with synthetic examples that were either created by Qwen2.5-Coder-32B Hui et al. (2024) or enhanced by Miqu-70B  [^17] https://huggingface.co/miqudev/miqu-1-70b. Every comment underwent automated validation against language-specific docstring conventions (GoogleDoc  [^18] https://google.github.io/styleguide/pyguide.html, JSDoc  [^19] https://jsdoc.app, etc.) using custom verification tools. The validation protocol enforced strict structural completeness, requiring all comments to comprehensively describe parameters, return values, exceptions, and types (if needed). For this benchmark, we curated a balanced subset comprising $7500$ comments for training ($1500$ per language) and $500$ for testing ($100$ per language).
For evaluation, the chrF metric Popović (2015) is employed due to its enhanced sensitivity to Russian morphological complexity. Unlike BLEU Papineni et al. (2002a) – which often fails to capture inflectional nuances – chrF operates through character n-gram weighting. This approach effectively detects subtle morphological shifts in case endings, verb conjugations, and derivational suffixes that are essential for assessing documentation fluency in Russian. The metric’s design aligns with the language’s rich inflectional system, providing a more reliable quality assessment for generated comments.
An example of StRuCom task:
  • instruction: Write the Russian-language documentation for the function.
  • inputs:
    private void button15_Click(object sender, EventArgs e) label12.Text = ""; richTextBox1.Clear(); richTextBox2.Clear(); textBox1.Clear(); textBox2.Clear(); textBox4.Clear();
  • outputs (golden answer):
    /// <summary>
    /// Handler for the button click event button15.
    /// When calling this function, the text in various controls of the form is cleared:
    /// - The text is deleted from the label label12.
    /// - The contents of the multiline text field richTextBox1 are cleared.
    /// - The contents of the multiline text field richTextBox2 are cleared.
    /// - The text field is cleared TextBox1.
    /// - The textBox2 text field is cleared.
    /// - The textbox textbox is being cleared./// </summary>
    /// <param name="sender">The object that triggered the event (in this case, the button15 button).
    </param>
    /// <param name="e">Event parameters that contain additional information about the event.</param>"
YABLoCo
The YABLoCo benchmark evaluates the proficiency of modern large language models (LLMs) in completing code functions based on textual descriptions from the user. While previous research has explored this capability for standalone functions Jiang et al. (2024), there has been limited investigation into functions intended for integration within larger code repositories. YABLoCo consists of 4 large repositories, with sizes ranging from 200,000 to 2,000,000 lines of code. The authors of the benchmark generated a call graph for each repository. For example, the LLVM repository contains approximately 175,000 functions and 571,000 inter-function calls. Next, the functions that do not correspond to the following criteria were filtered out from the benchmark: low test coverage, lack of developer comments, and excessively long function bodies. From the remaining candidates, the authors of YABLoCo manually selected 208 functions that featured clear docstrings aligned with their respective function bodies. These docstrings serve as the descriptive prompts for the LLMs being tested.
The main challenge for an LLM in generating a function for a specific code repository lies in the context size of the existing codebase. The new function must not only precisely follow the text description but also effectively utilize other functions and classes within the repository. This necessitates a careful selection of context that is both relevant for generation and not excessively lengthy. The authors of YABLoCo Valeev et al. (2025) discovered that an effective context for this purpose consists of the functions that are called by the target function. This context is referred to as the "oracle," as it is typically unavailable to the new function during generation, although it is known within the benchmark. Simply incorporating the oracle functions, without any modifications to the LLMs, resulted in an appreciable improvement on YABLoCo in the overall pass@10 metric, increasing it by 14 points (from 22.4 to 36.1).
An example of YABLoCo task:
  • instruction: Generate code on C.
  • inputs:
    Function signature:
    void *UI_add_user_data(UI *ui, void *user_data).
    Function description:
    The following function is used to store a pointer to user-specific data.
    Use this context:
    LONG_CONTEXT
  • outputs (golden answer):
    void *UI_add_user_data(UI *ui, void *user_data)
    void *old_data = ui->user_data;
    ui->user_data = user_data;
[…]
    return old_data;
Figure 2: The user path for the submission process on the MERA Code evaluation platform.
Appendix B Taxonomy details
In the Table 3 we provide an explanation of skills presented in the MERA Code taxonomy, starting from the second level of taxonomy.
Taxon Base Taxon Description Example of task
Text Perception Ability to take as input text modality MERA Code: StRuCom, UnitTests
Image Perception Ability to take as input image modality UML Diagrams, Block Schemes
Audio Perception Ability to take as input audio modality Vibe Coding
Video Perception Ability to take as input video modality Programming tutorial captioning
Code Perception Ability to take as input code files HTML, Markdown, Pseudo-code, bash scripts
Tools Perception Ability to take information from various tools Calculator, Web-Search, Interpreter
Long Context Comprehension Perception Ability to take as input sequences with length at least 32000 tokens.
Style Guides Knowledge Standards of code style for certain programming language MERA Code: CodeLinterEval(PEP8)
Programming Patterns Knowledge Programming patterns and best practices for developing a software project GoF
Programming Languages Knowledge Syntax of specific programming language Python, Java, C++
Information & Cyber Security Knowledge Awareness of model about threats in information systems Malware, hacking, phishing, ransomware
Algorithms & Data Structures Knowledge Theoretical knowledge in Computer Science Sorting algorithms, Asymptotic complexity
Comparison Reasoning Compare several code snippets with each other and derive their differences and commonalities Commit message diff
Side-by-Side Evaluation Reasoning Score a pair of code snippets according to certain criteria Code Review
Code-to-Code Search Reasoning Identify similar code snippets based on their syntactical and semantical similarity Duplicates identification
Knowledge Extraction Reasoning Find information in a given context relevant to a certain user query Question answering
Logs analysis Reasoning Extract particular events or summarize logs of any program execution Identify the reason of code crashing
Text-to-Code Search Reasoning Semantic search of the code snippet based on a given query Question answering
Dependencies understanding Reasoning Understanding of dependencies required to run certain application Requirements list creation
Classification Reasoning Classify the code snippet according to a specific labels MERA Code: Code correctness classification
Detection Reasoning Highlight and classify the span of code snippet Vulnerability detection
Segmentation Reasoning Classify every unit of input message(token, pixel, etc) to a specific labels OCR
Simulation Reasoning Simulation the behavior of code snippet after running MERA Code: StRuCom
Testing Reasoning Design a test suit for a program with the fully possible test coverage MERA Code: JavaTestGen
Review Reasoning Critical analysis of provided code snippet MERA Code: ruCodeReviewer
Instruction Following Reasoning Understanding of a provided instruction and adjustment of the answer according to it MERA Code: ruCodeReviewer
Planning Reasoning Plan future actions to achieve the certain goal End-to-End software development
Synthesis Generation Generation of code from scratch by a provided text prompt. MERA Code: Unit Tests generation
Completion Generation Generation of code based on provided code snippet and optional text instruction MERA Code: RealCode, RealCodeJava, ruCodeEval
Editing Generation Edit a provided code snippet accordingly to text instruction MERA Code: CodeLinterEval
Translation Generation Translate code from one programming language to another. Rewriting the code-base to a new programming language
Text Generation Generation Generation of text based on a ginen code Summarization of code snippet into natural language
Explanation Generation Textual description and explanation of code snippet README creation
Documentation Generation Generate structured explanation of code snippet MERA Code: StRuCom
Diagram Generation Generation Create visual description of software project UML Diagram creation
Table 3: The MERA Code taxonomy overview. It represents skills beginning from the second layer of taxonomy. Taxon - name of skill, Base Taxon - name of one of four grounding skills which is a predecessor of certain skill, Description - characteristic of skill, Example of task - real-world task where a mentioned skill could be used. Examples include tasks from MERA Code, if there are exists. MERA Team works on the extension of current benchmark to cover all skills, presented in the taxonomy.
Appendix C Evaluation results
This section provides two tables: Table 4 shows all metrics of the baseline models on private tasks of MERA Code, Table 5 provides the metrics for the same models on public tasks of the benchmark.
Model Private Score ruCodeEval CodeLinterEval ruCodeReviewer
pass@1
pass@5
pass@10
pass@1
pass@5
pass@10
BLEU
chrF
Judge@1
Judge@5
Judge@10
Gemini 2.5 flash 0.427 0.61 0.645 0.652 0.496 0.538 0.545 0.031 0.152 0.064 0.145 0.193
GPT-4.1 0.382 0.443 0.484 0.494 0.555 0.585 0.6 0.017 0.136 0.096 0.102 0.102
GPT-4o 0.381 0.529 0.559 0.567 0.479 0.518 0.527 0.018 0.133 0.078 0.094 0.1
GigaChat 2 Max 0.365 0.537 0.588 0.591 0.425 0.461 0.473 0.016 0.136 0.062 0.065 0.071
DeepSeek-Coder-V2-Inst 0.36 0.433 0.45 0.457 0.494 0.59 0.618 0.015 0.135 0.06 0.061 0.062
Seed-Coder-8B-Inst 0.345 0.317 0.317 0.317 0.655 0.655 0.655 0.019 0.156 0.035 0.052 0.06
Qwen2.5-Coder-32B-Inst 0.306 0.311 0.311 0.311 0.466 0.472 0.473 0.034 0.187 0.065 0.174 0.222
Qwen2.5-72B-Inst 0.254 0.174 0.177 0.177 0.481 0.497 0.5 0.023 0.158 0.048 0.104 0.136
Vikhr-YandexGPT-5-Lite-8B 0.187 0.035 0.041 0.043 0.407 0.515 0.518 0.015 0.126 0.022 0.023 0.023
Yi-Coder-9B-Chat 0.181 0.35 0.362 0.372 0.145 0.157 0.164 0.009 0.078 0.016 0.016 0.016
GigaCode 1.4 0.166 0.357 0.364 0.366 0.027 0.027 0.027 0.029 0.182 0.064 0.12 0.145
Mixtral-8x22B-Inst 0.028 0.0 0.0 0.0 0.027 0.045 0.045 0.017 0.135 0.016 0.025 0.025
Table 4: MERA Code benchmark results on private tasks. The best results are highlighted in bold.
Model Total Score
YABLoCo
stRuCom
RealCode
UnitTests
JavaTestGen
ruHumanEval
RealCodeJava
CodeCorrectness
pass@1
EM
chrF
pass@1
CBLEU
pass@1
compile@1
pass@1
pass@5
pass@10
pass@1
acc
GPT-4.1 0.377 0.144 0.034 0.297 0.418 0.162 0.344 0.639 0.45 0.48 0.494 0.416 0.66
GPT-4o 0.377 0.149 0.038 0.275 0.324 0.153 0.37 0.705 0.537 0.558 0.561 0.399 0.666
Gemini 2.5 flash 0.356 0.12 0.029 0.217 0.388 0.174 0.211 0.502 0.604 0.654 0.665 0.386 0.404
DeepSeek-Coder-V2-Inst 0.347 0.149 0.034 0.2 0.363 0.153 0.269 0.581 0.392 0.411 0.415 0.383 0.714
GigaChat 2 Max 0.346 0.106 0.014 0.294 0.332 0.175 0.137 0.396 0.53 0.57 0.579 0.342 0.68
Qwen2.5-Coder-32B-Inst 0.296 0.111 0.019 0.213 0.323 0.168 0.22 0.529 0.289 0.293 0.293 0.383 0.519
GigaCode 1.4 0.289 0.135 0.038 0.276 0.322 0.189 0.313 0.639 0.305 0.305 0.305 0.352 0.676
Qwen2.5-72B-Inst 0.285 0.144 0.024 0.252 0.362 0.128 0.189 0.476 0.157 0.163 0.165 0.349 0.702
Seed-Coder-8B-Inst 0.268 0.106 0.01 0.237 0.106 0.128 0.264 0.643 0.21 0.219 0.22 0.305 0.403
Yi-Coder-9B-Chat 0.203 0.135 0.024 0.192 0.067 0.138 0.229 0.59 0.173 0.197 0.201 0.252 0.364
Mixtral-8x22B-Inst 0.179 0.106 0.019 0.152 0.18 0.159 0.229 0.502 0.0 0.0 0.0 0.366 0.597
Vikhr-YandexGPT-5-Lite-8B 0.168 0.091 0.01 0.138 0.032 0.106 0.123 0.405 0.024 0.027 0.03 0.201 0.464
Table 5: MERA Code benchmark results on public tasks. The best results are highlighted in bold.